# Bytes to Schlep? Use a FEP: Hiding Protocol Metadata with Fully Encrypted Protocols

### v1

@@ 1. Introduction
One approach to avoid network censorship is to use a Fully Encrypted Protocol (FEP). FEPs are designed to hide communications metadata, such as the precise length of a plaintext message, the encryption algorithms being used, and the exact protocol being run. A goal of FEPs is that each byte appears uniformly random (a design sometimes referred to as “obfuscated”, “look-like-nothing”, “randomized”, or “unfingerprintable” (Dixon et al., 2016; Wu et al., 2023; Wang et al., 2015; Perrin, 2018)). Censorship circumvention tools using FEPs include the popular shadowsocks (sha, 2022), obfs4 (obfs4, 2024), and VMess (vmess, 2024) systems, each of which provides its own unique FEP. The Noise protocol framework (Perrin, 2018) requires ciphertexts be indistinguishable from random to be “censorship-resistant” (although it fails to specify all protocol aspects needed to satisfy that goal). While methods to identify and block FEPs exist (Wang et al., 2015; Wails et al., 2024), FEPs nonetheless continue to be effective and popular (sha, 2022; v2ray, 2024; obfs4, 2024).
@@
Despite these motivatations, FEPs have primarily been developed by the open-source community. Consequently, real-world protocols have been developed without mathematical security goals, even though their security depends on cryptographic assumptions [^1] The Pseudorandom cTLS proposal (Schwartz and Patton, 2022) cites the need for such formalization: “TODO: More precise security properties and security proof. The goal we’re after hasn’t been widely considered in the literature so far, at least as far as we can tell.”. Also, in practice, different developers make a variety of design errors (Fifield, 2023), as they lack a common set of FEP techniques and pitfalls.
Therefore, we propose novel, precise security definitions for FEPs. We formalize the goal of producing apparently random protocol messages with a passive adversary, and then we extend that to the goal of appearing otherwise predictable (and thus not leaking information) with respect to an active adversary. The predictability applies to the protocol’s behavior when ciphertexts are modified and to the protocol’s use of channel closures, an observable feature of connection-based transport protocols such as QUIC and TCP. Research has shown that channel closures can be used to identify specific FEPs (Frolov et al., 2020), and real-world measurements indicate that censors are already doing so (Beznazwy and Houmansadr, 2020).
In addition, most FEPs (and many other encrypted protocols) include message padding as a way to prevent traffic analysis from making inferences on traffic contents based on ciphertext lengths. For FEPs, an effective padding strategy is particularly powerful because each byte appears random, and so the number of bytes is the main remaining source of communications metadata. However, existing padding mechanisms are ad hoc and have no precisely stated goals. As a consequence, they fail to provide full control over over ciphertext lengths, a problem that has been observed by FEP developers themselves (Fifield, 2017), serves as the basis for state-of-the-art FEP detection using packet lengths (Wails et al., 2024), and has been exploited in practice by censors (Beznazwy and Houmansadr, 2020). We thus propose a powerful property called Traffic Shaping, which requires that protocols be capable of producing ciphertexts of arbitrary lengths on command.
@@
Finally, we analyze a wide variety of existing FEPs under our proposed security definitions, including the previously mentioned shadowsocks (sha, 2022), obfs4 (obfs4, 2024), and VMess (vmess, 2024) protocols. We examine their source code and documentation, and we run experiments to measure their error responses and message sizes. Our results indicate that nearly all of them do produce outputs that are indistinguishable from randomness, satisfying our passive FEP security definition. However, among the datastream protocols, their channel closures in response to errors make them identifiable, and they consequently fail our active FEP security definition. Our experiments further uncover integrity failures in V2Ray (the system providing the VMess protocol (v2ray, 2024)). Moreover, all of the FEPs exhibit unique minimum message sizes, which we both predict via source-code analysis and verify experimentally. The catalog we produce of channel closures and minimum message sizes provides a new method by which existing FEPs can be individually identified and thus blocked. It also supports our proposed security definitions, which normalize channel closures and preclude minimum message sizes.
@@ 2. Background and Preliminaries
We use a generic AEAD encryption scheme for our channel constructions in Sections 5.1 and 8.1, as well as certain counterexamples. The scheme consists of a triple of algorithms $\textsc{Gen}(\lambda)$, $\textsc{Enc}_{k}(\texttt{nonce},m,\texttt{ad})$, $\textsc{Dec}_{k}(\texttt{nonce},c,\texttt{ad})$, as defined by Rogaway (Rogaway, 2002), with both the AUTH and PRIV properties for integrity and confidentiality, respectively. We denote decryption errors with the distinguished error symbol $\bot_{\textsc{Dec}}$. We assume that the scheme satisfies IND\$-CPA (Rogaway, 2011), meaning that its ciphertext outputs are indistinguishable from random bytes. We further assume that the scheme has a fixed nonce size, $\ell_{\textrm{Nonce}}$, and assume that the scheme has a fixed overhead associated with encryption in the form of an additive constant, $\ell_{\textrm{Tag}}$, meaning that $|\textsc{Enc}_{k}(\texttt{nonce},m,\texttt{ad})|=|m|+\ell_{\textrm{Tag}}$ for any valid key $k$, nonce nonce, associated data ad, and message $m$. We note that this property implies length regularity; that is, input plaintexts of the same length produce output ciphertexts of the same length. Standard AEAD schemes, such as AES-GCM, are believed to satisfy all of the properties we assume.
@@ 3.2. Correctness
Let $S$ be a sequence of channel operations, starting with Init and followed by (possibly interleaved) calls to Send and Recv. Let $n_{s}$ be the number of Send calls and $n_{r}$ be the number of Recv calls. Let $M$, $P$, and $F$ be lists containing the inputs to the Send calls, $C$ be a list of the Send outputs, $C^{\prime}$ be a list of the Recv inputs, and $M^{\prime}$ and CL be lists of the Recv outputs. A channel satisfies correctness with respect to close function $\mathscr{C}$ if, for all such sequences $S$ where $\bigparallel C^{\prime}\preceq\bigparallel C$, the following properties hold:
@@ 3.3. Secure Close Functions
Finally, secure close functions are realistic. Both academic and real-world protocols have close behaviors which can, in whole or in part, be realized by a secure close function, such as the fixed-boundary closures of InterMAC (Albrecht et al., 2019) and the never-close behavior of shadowsocks in one direction (sha, 2022). We also observe a real-world attempt, in the VMess protocol (vmess, 2024), to obscure when ciphertext errors are detected by randomizing the timing of a subsequent closure, which doesn’t quite satisfy secure closures and consequently leaks metadata (see § 9).
@@ 3.5. Fully Encrypted Datastream Protocols
A channel satisfies FEP-$x$, for $x\in\{\textrm{CPFA},\textrm{CCFA}\}$ if, for any PPT adversary $\mathcal{A}$, $\left|P\left[\mathsf{Exp}^{\textsf{FEP-$x$},b}_{\mathcal{A}}(1^{\lambda})=1|b\overset{{\scriptscriptstyle\operatorname{R}}}{\leftarrow}\{0,1\}\right]-1/2\right|$ is negligible in the security parameter $\lambda$.
@@
2:$C_{S},C_{R},C_{\texttt{CL}}\leftarrow[],[],[]$
@@ 3.6. Traffic Shaping
A channel satisfies Traffic Shaping if, for any state $\texttt{st}_{S}$, message $m$, and integer $p\geq 0$, with $(\texttt{st}_{S},c_{f})\leftarrow\textsc{Send}(\texttt{st}_{S},m,p,f)$, $f\in\{0,1\}$, $|c_{0}|=p$ and $|c_{1}|\geq p$.
@@
Traffic shaping enables protection of protocol metadata by setting the $p$ inputs to values that are independent of the protocol being used. For example, the $p$ values could all be set to a constant, which would hide the number and sizes of metadata fields as well as the length of the plaintext messages. This is similar to the strategy used by the Tor protocol of putting all data into fixed-size cells to prevent traffic analysis (tor, 2023).
@@ 4. Relations Between Datastream Notions
CH-REGFEP-CCFAFEP-CPFAIND-CPFA-CLINT-CSTERR-PREDIND-CPFAIND-CCFAERR-FREEINT-PSTTheorem 4TrivialTrivialTheorem 2Theorem 3Theorem 5Trivial
@@
Theorem 3 shows that if a channel is correct and provides IND-CFPA (the standard confidentiality notion), then it provides the similar confidentiality notion with closures, IND-CPFA-CL.
@@ 5.1. The Construction
$\textsc{Init}(1^{\lambda})$: 1:$K\leftarrow_{\$}\mathcal{K}$ 2:$\texttt{st}_{S}=(K,0,\epsilon,\epsilon)$ 3:$\texttt{st}_{R}=(K,0,\epsilon,0)$ 4:return $(\texttt{st}_{S},\texttt{st}_{R})$   $\textsc{Send}(\texttt{st}_{S},m,p,f)$: 1:$(K,\texttt{seqno},\texttt{buf},\texttt{obuf})\leftarrow\texttt{st}_{S}$ 2: $\texttt{buf}\leftarrow\texttt{buf}\|m$ 3:if $(|\texttt{obuf}|\geq p)\land((f=0)\lor(\texttt{buf}=\epsilon))$ then 4:     $c\leftarrow\texttt{obuf}[1..\textsc{Max}(p,f*|\texttt{obuf}|)]$ 5:     $\texttt{obuf}\leftarrow\texttt{obuf}\%c$ 6:     return $(\texttt{st}_{S},c)$ 7:$o\leftarrow\textsc{Min}(|\texttt{buf}|,\texttt{il})$ 8:$\ell_{p}\leftarrow 0$ 9:$\ell_{c}\leftarrow|\textsc{Enc}(\texttt{seqno},0^{(2+o)})|$ 10:while $(|\texttt{obuf}|+\ell_{c}+\ell_{\texttt{len}}<p)\land(\ell_{c}\neq\texttt{ol})$ do 11:     $\ell_{p}\leftarrow\ell_{p}+1$ 12:     $\ell_{c}\leftarrow|\textsc{Enc}(\texttt{seqno},0^{(2+o+\ell_{p})})|$ 13:$c\leftarrow\textsc{Enc}(\texttt{seqno},\ell_{c})$ 14:$\texttt{seqno}\leftarrow\texttt{seqno}+1$ 15:if $\ell_{c}>0$ then 16:     $m^{\prime}\leftarrow\ell_{p}\|(0^{\ell_{p}})\|\texttt{buf}[1..o]$ 17:     $\texttt{buf}\leftarrow\texttt{buf}\%\texttt{buf}[1..o]$ 18:     $c\leftarrow c\|\textsc{Enc}(\texttt{seqno},m^{\prime})$ 19:     $\texttt{seqno}\leftarrow\texttt{seqno}+1$ 20:$\texttt{obuf}\leftarrow\texttt{obuf}\|c$ 21:return $\textsc{Send}(\texttt{st}_{S},\epsilon,p,f)$ $\textsc{Recv}(\texttt{st}_{R},c)$: 1:$(K,\texttt{seqno},\texttt{buf},\texttt{fail})\leftarrow\texttt{st}_{R}$ 2:if $\texttt{fail}=1$ then 3:     return $(\texttt{st}_{R},\epsilon,0)$ 4:$\texttt{buf}\leftarrow\texttt{buf}\|c$ 5:$m\leftarrow\epsilon$ 6:while $|\texttt{buf}|\geq\ell_{\texttt{len}}$ do 7:     $\ell_{c}\leftarrow\textsc{Dec}(\texttt{seqno},\texttt{buf}[1..\ell_{\texttt{len}}])$ 8:     if $\ell_{c}=\bot_{\textsc{Dec}}$ then 9:         $\texttt{fail}\leftarrow 1$ 10:         return $(\texttt{st}_{R},\epsilon,0)$ 11:     else if $\ell_{c}=0$ then 12:         $\texttt{buf}\leftarrow\texttt{buf}\%\texttt{buf}[1..\ell_{\texttt{len}}]$ 13:         $\texttt{seqno}\leftarrow\texttt{seqno}+1$ 14:     else if $|\texttt{buf}|\geq\ell_{\texttt{len}}+\ell_{c}$ then 15:         $c^{\prime}\leftarrow\texttt{buf}[\ell_{\texttt{len}}+1..\ell_{\texttt{len}}+\ell_{c}]$ 16:         $\texttt{buf}\leftarrow\texttt{buf}\%\texttt{buf}[1..1+\ell_{\texttt{len}}+\ell_{c}]$ 17:         $m^{\prime}\leftarrow\textsc{Dec}(\texttt{seqno}+1,c^{\prime})$ 18:         $\texttt{seqno}\leftarrow\texttt{seqno}+2$ 19:         if $m^{\prime}=\bot_{\textsc{Dec}}$ then 20:              $\texttt{fail}\leftarrow 1$ 21:              return $(\texttt{st}_{R},\epsilon,0)$           22:         $\ell_{p}\leftarrow\textsc{min}(m^{\prime}[1,2],|m^{\prime}|-2)$ 23:         $m\leftarrow m\|m^{\prime}[3+\ell_{p}..]$ 24:     else 25:         break       26:return $(\texttt{st}_{R},m,0)$
We give a construction that satisfies all our security notions for the secure close function $\mathcal{C}\equiv 0$, presented in Figure 2. The Send function contains an input buffer for plaintext and produces an output buffer of ciphertext blocks (adding blocks that only contain padding when necessary), outputting fragments of these blocks as requested by the caller. When Send is called, it returns an appropriate number of bytes from this output buffer if possible, and if not it first recursively extends the buffer by some new ciphertext block(s) before returning. The Recv function waits to receive enough ciphertext to contain a length block, and then it parses the length block and its (possibly empty) paired payload block before removing the pair from its internal buffer, stripping padding as each block is processed. Our construction does not produce channel closures or error symbols.
@@
We note that our construction generalizes in some ways. Our buffering and padding approach can be easily adapted for other FEP designs, such as obfs4 (obfs4, 2024) and InterMAC (Boldyreva et al., 2012), so that they satisfy Traffic Shaping. Further, although timing is outside of our model, an implementation could close the connection based on a timeout, where the timeout must depend only how much time has elapsed since the last ciphertext fragment was received.

### v2

@@ Abstract.
Cryptography; Network Protocols; Censorship Circumvention; Fully Encrypted Protocols
copyright: usgovjournalyear: 2024doi: 3658644.3690198conference: Proceedings of the 2024 ACM SIGSAC Conference on Computer and Communications Security; October 14–18, 2024; Salt Lake City, UT, USAbooktitle: Proceedings of the 2024 ACM SIGSAC Conference on Computer and Communications Security (CCS ’24), October 14–18, 2024, Salt Lake City, UT, USAdoi: 10.1145/3658644.3690198isbn: 979-8-4007-0636-3/24/10ccs: Security and privacy Symmetric cryptography and hash functionsccs: Networks Network protocol designccs: Networks Presentation protocolsccs: Networks Network privacy and anonymityccs: Networks Security protocolsccs: Security and privacy Privacy-preserving protocolsccs: Security and privacy Cryptanalysis and other attacks
@@ 1. Introduction
One approach to avoid network censorship is to use a Fully Encrypted Protocol (FEP). FEPs are designed to hide communications metadata, such as the precise length of a plaintext message, the encryption algorithms being used, and the exact protocol being run. A goal of FEPs is that each byte appears uniformly random (a design sometimes referred to as “obfuscated”, “look-like-nothing”, “randomized”, or “unfingerprintable” (Dixon et al., 2016; Wu et al., 2023; Wang et al., 2015; Perrin, 2018)). Censorship circumvention tools using FEPs include the popular Shadowsocks (Shadowsocks., 2022), obfs4 (obfs4, The obfourscator), and VMess (VMess., 2024) systems, each of which provides its own unique FEP. The Noise protocol framework (Perrin, 2018) requires ciphertexts be indistinguishable from random to be “censorship-resistant” (although it fails to specify all protocol aspects needed to satisfy that goal). While methods to identify and block FEPs exist (Wang et al., 2015; Wails et al., 2024), FEPs nonetheless continue to be effective and popular (Shadowsocks., 2022; Project V., 2024; obfs4, The obfourscator).
@@
Despite these motivations, FEPs have primarily been developed by the open-source community. Consequently, real-world protocols have been developed without mathematical security goals, even though their security depends on cryptographic assumptions [^1] The Pseudorandom cTLS proposal (Schwartz and Patton, 2022) cites the need for such formalization: “TODO: More precise security properties and security proof. The goal we’re after hasn’t been widely considered in the literature so far, at least as far as we can tell.”. Also, in practice, different developers make a variety of design errors (Fifield, 2023), as they lack a common set of FEP techniques and pitfalls.
Therefore, we propose novel, precise security definitions for FEPs. We formalize the goal of producing apparently random protocol messages with a passive adversary, and then we extend that to the goal of appearing otherwise predictable (and thus not leaking information) with respect to an active adversary. The predictability applies to the protocol’s behavior when ciphertexts are modified and to the protocol’s use of channel closures, an observable feature of connection-based transport protocols such as QUIC and TCP. Research has shown that channel closures can be used to identify specific FEPs (Frolov et al., 2020), and real-world measurements indicate that censors are already doing so (Alice et al., 2020).
In addition, most FEPs (and many other encrypted protocols) include message padding as a way to prevent traffic analysis from making inferences on traffic contents based on ciphertext lengths. For FEPs, an effective padding strategy is particularly powerful because each byte appears random, and so the number of bytes is the main remaining source of communications metadata. However, existing padding mechanisms are ad hoc and have no precisely stated goals. As a consequence, they fail to provide full control over over ciphertext lengths, a problem that has been observed by FEP developers themselves (Fifield, 2017), which serves as the basis for state-of-the-art FEP detection using packet lengths (Wails et al., 2024), and has been exploited in practice by censors (Alice et al., 2020). We thus propose a powerful property called Traffic Shaping, which requires that protocols be capable of producing ciphertexts of arbitrary lengths on command.
@@
Our theoretical results are inspired by observing persistent problems in real-world FEP deployments and our work provides concrete guidance for future FEP development. Our security definitions provide explicit security goals for FEPs to satisfy, and allow FEP maintainers to prove their protocols secure. Our Traffic Shaping notions give FEP designers a length obfuscation capability that is maximally flexible and practical to implement. In addition our definitions of secure close functions highlight the importance of carefully implementing connection closures, and explicitly provides a narrow set of possible close behaviors which do not reveal metadata unnecessarily.
Finally, we analyze a wide variety of existing FEPs under our proposed security definitions, including the previously mentioned Shadowsocks (Shadowsocks., 2022), obfs4 (obfs4, The obfourscator), and VMess (VMess., 2024) protocols. We examine their source code and documentation, and we run experiments to measure their error responses and message sizes. Our results indicate that nearly all of them do produce outputs that are indistinguishable from randomness, satisfying our passive FEP security definition. However, among the datastream protocols, their channel closures in response to errors make them identifiable, and they consequently fail our active FEP security definition. Our experiments further uncover integrity failures in V2Ray (the system providing the VMess protocol (Project V., 2024)). Moreover, all of the FEPs exhibit unique minimum message sizes, which we both predict via source-code analysis and verify experimentally. The catalog we produce of channel closures and minimum message sizes provides a new method by which existing FEPs can be individually identified and thus blocked. It also supports our proposed security definitions, which normalize channel closures and preclude minimum message sizes.
@@
This paper improves and expands on an early version of this work (Fenske and Johnson, 2023) by including a more developed notion of close functions, security definitions and a construction in the datagram setting, analyses of relations between the new security notions and existing ones, and an analysis of and experiments on existing FEPs. The appendices of this paper also contain proofs and other details.
@@ 2. Background and Preliminaries
We use a generic AEAD encryption scheme for our channel constructions in Sections 5.1 and 8.1, as well as certain counterexamples. The scheme consists of a triple of algorithms $\textsc{Gen}(\lambda)$, $\textsc{Enc}_{k}(\texttt{nonce},m,\texttt{ad})$, $\textsc{Dec}_{k}(\texttt{nonce},c,\texttt{ad})$, as defined by Rogaway (Rogaway, 2002), with both the AUTH and PRIV properties for integrity and confidentiality, respectively. In the remainder, we drop the associated data argument since it is not necessary for our work. We denote decryption errors with the distinguished error symbol $\bot_{\textsc{Dec}}$. We assume that the scheme satisfies IND\$-CPA (Rogaway, 2011), meaning that its ciphertext outputs are indistinguishable from random bytes. We further assume that the scheme has a fixed nonce size, $\ell_{\textrm{Nonce}}$, a fixed authentication tag size $\ell_{\textrm{Tag}}$ and assume that the scheme has a fixed overhead associated with encryption in the form of an additive constant, which we refer to as $\ell_{\textrm{Overhead}}$ meaning that $|\textsc{Enc}_{k}(\texttt{nonce},m)|=|m|+\ell_{\textrm{Overhead}}$ for any valid key $k$, nonce nonce, and message $m$. We call schemes with this property length additive. We note that this property implies length regularity; that is, input plaintexts of the same length produce output ciphertexts of the same length. Standard AEAD schemes, such as AES-GCM, are believed to satisfy all of the properties we assume.
@@ 3.2. Correctness
Let $S$ be a sequence of channel operations, starting with Init and followed by (possibly interleaved) calls to Send and Recv. Assume each Send or Recv call receives as input the state produced by its previous invocation (or the output of Init on the first such call). Let $n_{s}$ be the number of Send calls and $n_{r}$ be the number of Recv calls. Let $M$, $P$, and $F$ be lists containing the inputs to the Send calls, $C$ be a list of the Send outputs, $C^{\prime}$ be a list of the Recv inputs, and $M^{\prime}$ and CL be lists of the Recv outputs. A channel satisfies correctness with respect to close function $\mathscr{C}$ if, for all such sequences $S$ where $\bigparallel C^{\prime}\preceq\bigparallel C$, the following properties hold:
@@ 3.3. Secure Close Functions
Finally, secure close functions are realistic. Both academic and real-world protocols have close behaviors which can, in whole or in part, be realized by a secure close function, such as the fixed-boundary closures of InterMAC (Albrecht et al., 2019) and the never-close behavior of Shadowsocks in one direction (Shadowsocks., 2022). We also observe a real-world attempt, in the VMess protocol (VMess., 2024), to obscure when ciphertext errors are detected by randomizing the timing of a subsequent closure, which doesn’t quite satisfy secure closures and consequently leaks metadata (see § 9).
@@ 3.5. Fully Encrypted Datastream Protocols
A channel satisfies FEP-$x$, for $x\in\{\textrm{CPFA},\textrm{CCFA}\}$ if, for any PPT adversary $\mathcal{A}$, $\left|P\left[\mathsf{Exp}^{\textsf{FEP-$x$},b}_{\mathcal{A}}(1^{\lambda})=1\big{|}b\overset{{\scriptscriptstyle\operatorname{R}}}{\leftarrow}\{0,1\}\right]-1/2\right|$ is negligible in the security parameter $\lambda$.
@@
2:$C_{S},C_{R},C_{\texttt{CL}},\texttt{sync}\leftarrow[],[],[],1$
@@ 3.6. Traffic Shaping
A channel satisfies Traffic Shaping if, for any state $\texttt{st}_{S}$, message $m$, and integer $p\geq 0$, with $(\texttt{st}_{S},c)\leftarrow\textsc{Send}(\texttt{st}_{S},m,p,f)$, if $f=0$ then $|c|=p$, otherwise $|c|\geq p$.
@@
Traffic shaping enables protection of protocol metadata by setting the $p$ inputs to values that are independent of the protocol being used. For example, the $p$ values could all be set to a constant, which would hide the number and sizes of metadata fields as well as the length of the plaintext messages. This is similar to the strategy used by the Tor protocol of putting all data into fixed-size cells to prevent traffic analysis (Dingledine and Mathewson, 2023).
@@ 4. Relations Between Datastream Notions
CH-REGFEP-CCFAFEP-CPFAIND-CPFA-CLINT-CSTERR-PREDIND-CPFAIND-CCFAERR-FREEINT-PSTThm 4TrivialTrivialThm 2Thm 3Thm 5Trivial
@@
Theorem 3 shows that if a channel is correct and provides IND-CPFA (the standard confidentiality notion), then it provides the similar confidentiality notion with closures, IND-CPFA-CL.
@@ 5.1. The Construction
[…]
We give a construction that satisfies all our security notions for the secure close function $\mathcal{C}\equiv 0$, presented in Figure 2. Our approach is inspired by Shadowsocks, and is designed around producing pairs of ciphertexts, the first of which is a “length block” which has a fixed length, denoted $\ell_{\texttt{len}}=2+\ell_{\textrm{Tag}}$, and contains the length of the subsequent “payload block”, which is limited to a length that can be represented within two bytes. The payload block also contains a two byte padding length field to denote internal (plaintext) padding. We denote the largest payload block length as $\texttt{ol}=2^{16}-1$ and set il as the largest plaintext length that can fit within a payload block ($2^{16}-3-\ell_{\textrm{Tag}}$).
[…]
The Recv function keeps a buffer of received ciphertext bytes and waits to receive $\ell_{\texttt{len}}$ bytes, then decrypts and parses the length block to recover the length of the subsequent payload block. It then waits until the full payload block has been received before decrypting and returning the plaintext, stripping off the padding length and any padding bytes. Recv checks for decryption errors and enters a fail state if either block type fails to decrypt. Our construction does not produce channel closures or error symbols.
@@
We note that our construction generalizes in some ways. Our buffering and padding approach can be easily adapted for other FEP designs, such as obfs4 (obfs4, The obfourscator) and InterMAC (Boldyreva et al., 2012), so that they satisfy Traffic Shaping. Further, although timing is outside of our model, an implementation could close the connection based on a timeout, where the timeout must depend only on how much time has elapsed since the last ciphertext fragment was received.